# Identification and estimation of mediational effects of longitudinal modified treatment policies

### v3

@@ 3.2 Inverse probability weighted identification
as well as $\mathsf{K}_{l,u}^{\prime}=\prod_{r=l}^{u}\mathsf{G}_{A,r}^{\prime}$, $\mathsf{K}_{l,u}^{\star}=\prod_{r=l}^{u}\mathsf{G}_{A,r}^{\star}$, $\mathsf{H}_{l,u}=\prod_{r=l}^{u}\mathsf{G}_{M,r}$. (For notational brevity, we have omitted the dependence on $A_{t},$, $H_{A,t}$, $H_{M,t}$, and $m_{t}$.) Then, the parameters from Theorem 1 can also be identified as $\varphi(\bar{m})=\mathsf{E}\{\mathsf{K}_{1,\tau}^{\prime}\mathsf{H}_{1,\tau}Y\}$; $\lambda(\bar{m})=\mathsf{E}\{\mathsf{K}_{1,\tau}^{\star}\mathds{1}\{\bar{M}=\bar{m}\}\}$; see Section S3 of the supplement for the derivation.
@@ 3.3 Overview of estimation
The foregoing analysis allows us to construct estimators of the parameter $\theta$ given in Theorem 1 with desirable statistical properties such as asymptotic normality and efficiency. We have described sequential regression-based and IPW-based strategies for clarity, but see Section S5 of the supplement for a description of how we leverage semiparametric theory to use both the outcome and treatment models to yield a combined estimator which is consistent if either one is correctly specified at each timepoint, and which can achieve $\sqrt{n}$-rate convergence even if the individual estimators converge at a slower rate. Our estimation algorithm has been implemented in the open-source R package lcmmtp (Williams and Díaz, 2023).
@@ 4.1 Background
Acute respiratory distress syndrome (ARDS) is a major cause of morbidity and mortality among COVID-19 patients. Treating ARDS often requires the use of respiratory support devices, ranging from nasal cannulae and face masks to more invasive methods such as mechanical ventilation via endotracheal tubes (“invasive mechanical ventilation” or IMV) (Hasan et al., 2020). Given the physiological links between the lungs and kidneys, it has been suggested that mechanical ventilation may cause acute kidney injury/failure (AKI); possible mechanisms include oxygen toxicity and capillary endothelial damage leading to inflammation, hypotension, and sepsis (Durdevic et al., 2020). AKI, which can render kidneys incapable of appropriately clearing critical toxins from the body or maintaining appropriate blood and interstitial volume, complicates about 30% of ICU admissions and increases the risk of in-hospital mortality and long-term morbidity and mortality (Kes and Jukić, 2010). While IMV may often be necessary to support respiration even if it can lead to AKI (Husain-Syed et al., 2016), the decision whether and when to intubate must be based on clinical factors such as oxygen saturation, dyspnea, respiratory rate, and chest radiography, weighing in the benefit of reducing ARDS-related mortality with the possibility of harm through AKI (Tobin, 2020; Thomson and Calligaro, 2021).
@@
This data was previously described and studied in Díaz et al. (2023a) and Hoffman et al. (2024). The latter considered only effects on mortality without regard to AKI. The former treated AKI and other-cause mortality as competing risks; we take a different perspective by considering AKI as a mediator in the pathway from IMV to mortality. This orientation presents a more intricate view of the causal connections among these variables. Essentially, AKI and mortality are not only competing risks; they can jointly occur, with AKI potentially influencing mortality. By examining AKI’s mediating role, we aim to unravel the indirect ways in which IMV may contribute to mortality via AKI. Should this causal pathway exist, it could provide novel opportunities for medical interventions aimed at preventing or treating AKI, subsequently reducing the mortality rate. (As mentioned in Section 1, these options include the modulation of intravascular volume status (Grams et al., 2011; Glassford and Bellomo, 2011) and various pharmacological interventions (Pickkers et al., 2022).) Further, we analyze how various baseline variables are associated with the estimated direct and indirect effects and discuss how such considerations might inform treatment decisions regarding IMV.
@@ 4.2 Time-varying variables, mediator, and modified treatment policy
This study uses data at the daily level, and study time begins on the day of hospitalization. The treatment variable is categorized into three levels: no supplemental oxygen, supplemental oxygen not including IMV, and IMV. The main goal is to estimate the overall causal effect of IMV on mortality/survival rates and to decompose it into effects operating through AKI and effects operating independently of it. To measure the total effect of IMV on survival, we posit an intervention that would delay IMV by one day among patients who received IMV for more than one day and would prevent IMV for patients who received IMV for exactly one day. We refer to this intervention as a “one-day delay in intubation.” This relatively modest intervention (as opposed to one that delayed intubation by long) is motivated by a desire to preserve positivity (2(i)) and avoid extreme inverse propensity score weights. In actuality, 381 of the patients received intubation during the time period considered.
@@
Formally, for $A_{t}\in\{0,1,2\}$ corresponding to the end of day $t$, 00 indicates no supplemental oxygen, $1$ indicates supplemental oxygen not including IMV, and $2$ indicates IMV. We consider the following intervention:
$\mathsf{d}_{t}(a_{t},h_{a,t})=\begin{cases}1&\text{ if }a_{t}=2\text{ and }a_{s}\leq 1\text{ for all }s<t,\\ a_{t}&\text{ otherwise.}\end{cases}$ (5)
@@
In words, this means that the intervention leaves the observed treatment value unchanged, except for the first day the patient receives IMV (if that day occurs). For that day, the intervention changes the treatment to supplemental oxygen without IMV. This operationalizes the idea of postponing IMV debated early in the pandemic.
@@
In constructing the analytical dataset, it is important that the variables preserve the time-ordering of the DAG in Figure 1. Therefore, we constructed the dataset using the following procedure. First, we categorize patients into three groups: those whose first event was an intubation, those whose first event was AKI, and those who had neither event. For patients in the first group, we anchor the creation of their record on the time at which the intubation event occurred, denoted $t_{I}$. The variables $L_{t_{I-1}}$ and $Z_{t_{I-1}}$ are then recorded as those observed in approximately 24-hour windows before $t_{I}$, where $L_{t_{I-1}}$ contains variables measured in the first half of this approximate 24-hour interval, and $Z_{t_{I-1}}$ contains variables measured in the second half of this approximate 24-hour window. The time interval is approximate because the difference in hours between hospital admission and intubation is not necessarily a multiple of 24. The mediator $M_{t_{I-1}}$ is then set as 00, and the exposure $A_{t_{I-1}}$ is recorded as the respiratory support in the previous 24-hour window. This process is iterated until we have data for all the timepoints prior to $t_{I}$. For times greater than or equal to $t_{I}$, we follow a similar approach, but in addition to capturing $Z_{t}$, $L_{t}$, and $A_{t}$, we measure whether the patient was diagnosed with AKI at every time window. AKI was defined using creatinine values in accordance with the “Kidney Disease: Improving Global Outcomes” definition (KDIGO) definition (Khwaja, 2012). Either of the following criteria was required: (a) serum creatinine change of greater than or equal to 0.3 mg/dL within 48 h, or (b) serum creatinine greater than or equal to 1.5 times the baseline serum creatinine known or assumed to have occurred within the past 7 days.
@@
If and when a patient is diagnosed with AKI (call this time $t_{M}$), we begin a new 24-hour interval, i.e. $L_{t_{M}}$, to ensure that the temporality of variables measured after $t_{M}$ is maintained. A similar procedure was used for patients whose first event was the diagnosis of AKI. Data for patients in the third group, who had neither event, was simply divided into approximate 24-hour time intervals using the anchor times of hospital admission and censoring or death date. During this process, we ensured that measurement of $L_{t}$ precedes $Z_{t}$ within each time interval. A GitHub repository of analysis code is available at https://github.com/CI-NYC/lcmmtp-application. For more details, including the list of measured covariates, see Section S6 of the supplement.
@@ 4.3 Results
We used our proposed methods to estimate the total, direct, and indirect effects on mortality of invasive mechanical ventilation through acute kidney failure. All regressions were estimated using lcmmtp (Williams and Díaz, 2023) and Super Learner (van der Laan et al., 2007), building a convex combination of three predictors: generalized linear models with $\ell_{1}$ penalty, multivariate adaptive regression splines, and extreme gradient boosting of regression trees. To speed up computations, we used three folds for cross-fitting and used four-day mortality as our outcome of interest. Some observations had large weights $\mathsf{K}_{t,s}$ and $\mathsf{H}_{t,s}$, so we truncated the corresponding estimates at their 99% quantiles to avoid empirical violations of the positivity assumption. The results of the main analysis are presented in Table LABEL:tab:total, where we contrast the mortality under an intervention to prevent censoring versus the mortality under an intervention to prevent censoring and to delay intubation by one day as specified above. Specifically, we estimate the total effect defined as $\mathsf{E}[Y(\bar{A}^{\mathsf{d}},\bar{J}(\bar{A}^{\mathsf{d}}))-Y(\bar{A},\bar{J}(\bar{A}))]$, where $\mathsf{d}$ is defined in (5) and $Y$ is four-day survival rate under an intervention to prevent censoring. The indirect effect is thus defined as $\mathsf{E}[Y(\bar{A},\bar{J}(\bar{A}^{\mathsf{d}}))-Y(\bar{A},\bar{J}(\bar{A}))]$, and the direct effect is $\mathsf{E}[Y(\bar{A}^{\mathsf{d}},\bar{J}(\bar{A}^{\mathsf{d}}))-Y(\bar{A},\bar{J}(\bar{A}^{\mathsf{d}}))]$.
Although there is high uncertainty in the results, the point estimates support the hypothesis that invasive mechanical ventilation reduces mortality through mechanisms other than acute kidney failure. However, the effect through acute kidney failure is larger than the effect through all other mechanisms, resulting in an overall harmful effect of invasive mechanical ventilation on mortality. The estimated value of the total effect indicates that delaying intubation by one day would reduce mortality by 5.9 percentage points, with a standard error of 3.8 percentage points.
@@ 4.4 Mediation and treatment decisions
By contrast, one could attempt to estimate the total effect for all patients and exclude any patients with a negative value, but this shuts out the possibility of additional decision-making input from physicians and ignores the possibility of ameliorating the indirect effects (i.e., that additional interventions could be put in place to mitigate the probability of AKI for patients under intubation).
@@ 6 Acknowledgments
Supplementary Materials for
Identification and estimation of mediational effects of longitudinal modified treatment policies
S1 Technical note on notation
In the proofs that follow, any variable with index $t\leq 0$ should be interpreted as null, and a distribution conditional on null variables are marginal. Any summation expression with no terms (e.g., $\sum_{t+1}^{t}W_{t}$) are equal to zero, and any product with no factors is equal to one.
S2 Identification (Theorem 1)
@@
$=\mathsf{E}[\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))\mid J(\bar{A}^{\mathsf{d}^{\star}})]]$
$=\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))\mid\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m}]*\mathsf{P}(\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$
$=\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{m})\mid\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m}]*\mathsf{P}(\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$
@@
$=\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{m})]*\mathsf{P}(\bar{M}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$
@@
First, for $\varphi$, we consider the relabeled dataset where for all $t\leq\tau$, $(L_{t},Z_{t})=(C_{2t-1},C_{2t})$ and $(A_{t},M_{t})=(E_{2t-1},E_{2t})$. In this case, the expression $Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{m})$ is the expected value of the counterfactual outcome of the longitudinal modified treatment policy setting the values of $E$ to the corresponding values $\bar{A}^{\mathsf{d}^{\prime}}$ and $\bar{m}$. Similarly to Díaz et al. (2023b), this expression is identified by the following strategy.
@@
where $H_{t}$ denotes all variables prior to $E_{t}$. Then by integration, $\mathsf{E}[Y(\mathsf{E}^{\mathsf{d}^{\prime}})]=\mathsf{E}[\mu_{1}(E_{1}^{\prime},C_{1})]$.
Noting that the desired intervention on $\bar{m}$ is not stochastic (and thus the mediation component of the intervention $\mathsf{d}^{\prime}$ can be passed into the $\mathsf{Q}$-functions as the future histories $\underaccent{\bar}{m}_{t}$); thus one can see that substituting variables is sufficient to recognize, when $\mathsf{d}^{\prime}$ is the intervention setting $E$ to the values of $\bar{A}^{\mathsf{d}^{\prime}}$ and $\bar{m}$, the relation $\mathsf{E}[\mu_{1}(E_{1}^{\prime},C_{1})]=\mathsf{Q}_{L,0}^{\prime}(\bar{m})$, as in the Theorem.
For $\lambda(\bar{m})=\mathsf{P}(\bar{M}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$, we again relabel the data, this time grouping $L,Z,M$ as $S_{t}=(L_{t},Z_{t-1},\bar{M}_{t-1})$ (and again suitably altering the meaning of $H_{t}$ to denote all variables before $A_{t}$, or $H_{A,t}$ in the original labeling). Then $\bar{M}(\bar{A}^{\mathsf{d}\star})$ is simply a projection of the outcome $S_{\tau+1}$ under the longitudinal modified treatment policy $\bar{A}^{\mathsf{d}\star}$. Denoting this projection as $S_{t}^{m}=\bar{M}_{t-1}$, we can write $\mathsf{P}(\bar{M}(\bar{A}^{\mathsf{d}\star})=\bar{m})=\mathsf{E}[\mathbb{1}(S_{\tau+1}^{m}=\bar{m})]$ and we can again apply a similar identification strategy.
As a shorthand, let $\mathbb{1}_{\tau+1}^{\bar{m}}=\mathbb{1}(S_{\tau+1}^{m}=\bar{m})$. Let $\rho_{\tau+1}=\mathbb{1}_{\tau+1}^{\bar{m}}$ and for $t=\tau,\dots,1$ recursively define
@@
$=\mathbb{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{P}(M_{\tau}=m_{\tau}\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau})$
$=\mathbb{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{E}[\mathbb{1}(M_{\tau}=m_{\tau})\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau}]$
@@
$=\mathbb{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{Q}_{M,\tau}^{*}(\underaccent{\bar}{m}_{\tau})$
@@
$=\mathsf{E}[\mathbb{1}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{Q}_{M,\tau}^{*}(\underaccent{\bar}{m}_{\tau})\mid A_{\tau-1}=a_{\tau-1},H_{t}=h_{\tau-1}]$
$=\mathbb{1}(S_{\tau-1}^{m}=\bar{m}_{\tau-2})\mathsf{E}[\mathbb{1}(M_{\tau-1}=m_{\tau-1})\mathsf{Q}_{M,\tau}^{*}(\underaccent{\bar}{m}_{\tau})\mid A_{\tau-1}=a_{\tau-1},H_{t}=h_{\tau-1}]$
$=\mathbb{1}(S_{\tau-1}^{m}=\bar{m}_{\tau-2})\mathsf{Q}_{M,\tau-1}^{\star}(\underaccent{\bar}{m}_{\tau-1})$
@@
$\rho_{t}(a_{t},h_{t})=\mathbb{1}(S_{t}^{m}=\bar{m}_{t-1})\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})$
@@
S3 IPW identification
Díaz et al. (2021) gives, for a general longitudinal modified treatment policy $\bar{A}^{\mathsf{d}}$ and outcome $Y$ (without explicit mediation or intermediate confounders):
@@
S4 More on interpretation of interventional Effects
@@
Interventional direct and indirect effects offer an alternative framework that overcomes this limitation. The interventional direct effect (IDE) is defined as the difference in expected outcomes when the exposure is set to $a$ and the mediator is intervened to follow the distribution it would have under exposure $a^{\prime}$, compared to when both the exposure and mediator are set to $a^{\prime}$. Mathematically, this can be expressed as:
@@
which is not necessarily equal to $\mathsf{E}[Y(a)]-\mathsf{E}[Y(a^{\prime})]$. This is in contrast to the natural effects, since $\mathsf{E}[Y(a,M(a))]-\mathsf{E}[Y(a^{\prime},M(a^{\prime}))]$ does equal $\mathsf{E}[Y(a)]-\mathsf{E}[Y(a^{\prime})]$ since $Y(a,M(a))=Y(a)$).

### v4

@@ 3.2 Inverse probability weighted identification
as well as $\mathsf{K}_{l,u}^{\prime}=\prod_{r=l}^{u}\mathsf{G}_{A,r}^{\prime}$, $\mathsf{K}_{l,u}^{\star}=\prod_{r=l}^{u}\mathsf{G}_{A,r}^{\star}$, $\mathsf{H}_{l,u}=\prod_{r=l}^{u}\mathsf{G}_{M,r}$. (For notational brevity, we have omitted the dependence on $A_{t}$, $H_{A,t}$, $H_{M,t}$, and $m_{t}$.) Then, the parameters from Theorem 1 can also be identified as $\varphi(\bar{m})=\mathsf{E}\{\mathsf{K}_{1,\tau}^{\prime}\mathsf{H}_{1,\tau}Y\}$; $\lambda(\bar{m})=\mathsf{E}\{\mathsf{K}_{1,\tau}^{\star}\mathds{1}\{\bar{M}=\bar{m}\}\}$; see Section S3 of the supplement for the derivation.
@@ 3.3 Overview of estimation
The foregoing analysis allows us to construct estimators of the parameter $\theta$ given in Theorem 1. When using machine learning for estimation of the outcome and treatment mechanisms of the previous section, it is desirable to use doubly robust estimators, which guarantee efficiency and asymptotic normality under certain assumptions (Kennedy, 2024). For our problem, the construction of those estimators follows from a straightforward generalization of the methods of D´ıaz et al. (2021) and D´ıaz et al. (2023b). These estimators leverage semiparametric theory to use both the outcome and treatment models to yield a combined estimator which is consistent if either one is correctly specified at each timepoint, and which can achieve $\sqrt{n}$-rate convergence even if the individual estimators converge at a slower rate. Please see Section S5 of the supplement for a complete description of the estimators. Our estimation algorithm has been implemented in the open-source R package lcmmtp (Williams and D´ıaz, 2023).
@@ 4.1 Background
Acute respiratory distress syndrome (ARDS) is a major cause of morbidity and mortality among COVID-19 patients. Treating ARDS often requires the use of respiratory support devices, ranging from nasal cannulae and face masks to more invasive methods such as mechanical ventilation via endotracheal tubes (“invasive mechanical ventilation” or IMV) (Hasan et al., 2020). Given the physiological links between the lungs and kidneys, it has been suggested that mechanical ventilation may cause acute kidney injury/failure (AKI); possible mechanisms include oxygen toxicity and capillary endothelial damage leading to inflammation, hypotension, and sepsis (Durdevic et al., 2020). AKI, which can render kidneys incapable of appropriately clearing critical toxins from the body or maintaining appropriate blood and interstitial volume, complicates about 30% of ICU admissions and increases the risk of in-hospital mortality and long-term morbidity and mortality (Kes and Jukić, 2010).
@@
This data was previously described and studied in D´ıaz et al. (2023a) and Hoffman et al. (2024). The latter considered only effects on mortality without regard to AKI. The former treated AKI and other-cause mortality as competing risks; we take a different perspective by considering AKI as a mediator in the pathway from IMV to mortality. This orientation presents a more intricate view of the causal connections among these variables. Essentially, AKI and mortality are not only competing risks; they can jointly occur, with AKI potentially influencing mortality. By examining AKI’s mediating role, we aim to unravel the indirect ways in which IMV may contribute to mortality via AKI. Should this causal pathway exist, it could provide novel opportunities for medical interventions aimed at preventing or treating AKI, subsequently reducing the mortality rate. (As mentioned in Section 1, these options include the modulation of intravascular volume status (Grams et al., 2011; Glassford and Bellomo, 2011) and various pharmacological interventions (Pickkers et al., 2022).) Further, we analyze how various baseline variables are associated with the estimated direct and indirect effects and discuss how such considerations might inform treatment decisions regarding IMV.
@@ 4.2 Time-varying variables, mediator, and modified treatment policy
This study uses data at the daily level, and study time begins on the day of hospitalization. The treatment variable is categorized into three levels: no supplemental oxygen, supplemental oxygen not including IMV, and IMV. The main goal is to estimate the overall causal effect of IMV on mortality/survival rates and to decompose it into effects operating through AKI and effects operating independently of it. To measure the total effect of IMV on survival, we posit an intervention that would assign a noninvasive type of oxygen support to patients requiring oxygen support. This relatively modest intervention (as opposed to one that requires everyone to receive the same level of oxygen support) is motivated by a desire to preserve positivity (2,(i)) and avoid extreme inverse propensity score weights. In actuality, 381 of the patients received intubation during the time period considered.
@@
Formally, for $A_{t}\in\{0,1,2\}$ corresponding to the end of day $t$, 0 indicates no supplemental oxygen, $1$ indicates supplemental oxygen not including IMV, and $2$ indicates IMV. We consider the following intervention:
$\mathsf{d}_{t}(a_{t})=\begin{cases}1&\text{ if }a_{t}=2\\ a_{t}&\text{ otherwise.}\end{cases}$ (5)
@@
In words, this means that the intervention leaves the observed treatment value unchanged, except for days when the patient was intubated. For those days, the intervention changes the treatment to supplemental oxygen without IMV.
@@
In constructing the analytical dataset, it is important that the variables preserve the time-ordering of the DAG in Figure 1. Therefore, we constructed the dataset using the following procedure. First, we categorize patients into three groups: those whose first event was an intubation, those whose first event was AKI, and those who had neither event. For patients in the first group, we anchor the creation of their record at the time at which the intubation event occurred, denoted $t_{I}$. The variables $L_{t_{I-1}}$ and $Z_{t_{I-1}}$ are then recorded as those observed in approximately 24-hour windows before $t_{I}$, where $L_{t_{I-1}}$ contains variables measured in the first half of this approximate 24-hour interval, and $Z_{t_{I-1}}$ contains variables measured in the second half of this approximate 24-hour window. The time interval is approximate because the difference in hours between hospital admission and intubation is not necessarily a multiple of 24. The mediator $M_{t_{I-1}}$ is then set as 0, and the exposure $A_{t_{I-1}}$ is recorded as the respiratory support in the previous 24-hour window. This process is iterated until we have data for all the timepoints prior to $t_{I}$. For times greater than or equal to $t_{I}$, we follow a similar approach, but in addition to capturing $Z_{t}$, $L_{t}$, and $A_{t}$, we measure whether the patient was diagnosed with AKI at every time window. AKI was defined using creatinine values in accordance with the “Kidney Disease: Improving Global Outcomes” (KDIGO) definition (Khwaja, 2012). Either of the following criteria was required: (a) serum creatinine change of greater than or equal to 0.3 mg/dL within 48 h, or (b) serum creatinine greater than or equal to 1.5 times the baseline serum creatinine known or assumed to have occurred within the past 7 days.
@@
If and when a patient is diagnosed with AKI (call this time $t_{M}$), we begin a new 24-hour interval, i.e., $L_{t_{M}}$, to ensure that the temporality of variables measured after $t_{M}$ is maintained. A similar procedure was used for patients whose first event was the diagnosis of AKI. Data for patients in the third group, who had neither event, was simply divided into approximate 24-hour time intervals using the anchor times of hospital admission and censoring or death date. During this process, we ensured that measurement of $L_{t}$ precedes $Z_{t}$ within each time interval. A GitHub repository of analysis code is available at https://github.com/CI-NYC/lcmmtp-application. For more details, including the list of measured covariates, see Section S6 of the supplement.
@@ 4.3 Results
We used our proposed methods to estimate the total, direct, and indirect effects on mortality of invasive mechanical ventilation through acute kidney injury. All regressions were estimated using lcmmtp (Williams and D´ıaz, 2023) and Super Learner (van der Laan et al., 2007), building a convex combination of three predictors: generalized linear models with $\ell_{1}$ penalty, multivariate adaptive regression splines, and extreme gradient boosting of regression trees. To speed up computations, we used three folds for cross-fitting and used four-day survival as our outcome of interest. Some observations had large weights $\mathsf{K}_{t,s}$ and $\mathsf{H}_{t,s}$, so we truncated the corresponding estimates at their 99th percentiles to avoid empirical violations of the positivity assumption. The results of the main analysis are presented in Table LABEL:tab:total, where we contrast the mortality under an intervention to prevent censoring versus the mortality under an intervention to prevent censoring and to prevent intubation as discussed above. Specifically, we estimate the total effect defined as $\mathsf{E}[Y(\bar{A}^{\mathsf{d}},\bar{J}(\bar{A}^{\mathsf{d}}))-Y(\bar{A},\bar{J}(\bar{A}))]$, where $\mathsf{d}$ is defined in (5) and $Y$ is four-day survival rate under an intervention to prevent censoring. The indirect effect is thus defined as $\mathsf{E}[Y(\bar{A},\bar{J}(\bar{A}^{\mathsf{d}}))-Y(\bar{A},\bar{J}(\bar{A}))]$, and the direct effect is $\mathsf{E}[Y(\bar{A}^{\mathsf{d}},\bar{J}(\bar{A}^{\mathsf{d}}))-Y(\bar{A},\bar{J}(\bar{A}^{\mathsf{d}}))]$.
Although there is high uncertainty in the results, the point estimates support the hypothesis that invasive mechanical ventilation reduces mortality through mechanisms other than acute kidney injury. However, the effect through acute kidney injury is larger than the effect through all other mechanisms, resulting in an overall harmful effect of invasive mechanical ventilation on mortality. The estimated value of the total effect indicates that preventing intubation would reduce mortality by 5.9 percentage points, with a standard error of 3.8 percentage points.
@@ 4.4 Mediation and treatment decisions
By contrast, one could attempt to estimate the total effect for all patients and exclude any patients with a negative value, but this precludes the possibility of additional decision-making input from physicians and ignores the possibility of ameliorating the indirect effects (i.e., that additional interventions could be put in place to mitigate the probability of AKI for patients under intubation).
@@ 5 Discussion
6 Acknowledgments
Supplementary Materials
Appendix S1 Technical note on notation
In the proofs that follow, any variable with index $t\leq 0$ should be interpreted as null, and a distribution conditional on null variables is marginal. Any summation expression with no terms (e.g., $\sum_{t+1}^{t}W_{t}$) is equal to zero, and any product with no factors is equal to one.
Appendix S2 Identification (Theorem 1)
@@ 6 Acknowledgments
$\mathsf{E}[\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))\mid J(\bar{A}^{\mathsf{d}^{\star}})]]$
$\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{J}(\bar{A}^{\mathsf{d}^{\star}}))\mid\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m}]\cdot\mathsf{P}(\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$
$\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{m})\mid\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m}]\cdot\mathsf{P}(\bar{J}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$
@@
$\sum_{\bar{m}}\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{m})]\cdot\mathsf{P}(\bar{M}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$
@@
First, for $\varphi$, we consider the relabeled dataset where for all $t\leq\tau$, $(L_{t},Z_{t})=(C_{2t-1},C_{2t})$ and $(A_{t},M_{t})=(E_{2t-1},E_{2t})$. In this case, the expression $\mathsf{E}[Y(\bar{A}^{\mathsf{d}^{\prime}},\bar{m})]$ is the expected value of the counterfactual outcome of the longitudinal modified treatment policy setting the values of $E$ to the corresponding values $\bar{A}^{\mathsf{d}^{\prime}}$ and $\bar{m}$. Similarly to D´ıaz et al. (2023b), this expression is identified by the following strategy.
@@
where $H_{t}$ denotes all variables prior to $E_{t}$. Then by integration, $\mathsf{E}[Y(E^{\mathsf{d}^{\prime}})]=\mathsf{E}[\mu_{1}(E_{1}^{\prime},C_{1})]$.
Since the desired intervention on $\bar{m}$ is not stochastic (and thus the mediation component of the intervention $\mathsf{d}^{\prime}$ can be passed into the $\mathsf{Q}$-functions as the future histories $\underaccent{\bar}{m}_{t}$), it follows that substituting variables is sufficient to recognize, when $\mathsf{d}^{\prime}$ is the intervention setting $E$ to the values of $\bar{A}^{\mathsf{d}^{\prime}}$ and $\bar{m}$, the relation $\mathsf{E}[\mu_{1}(E_{1}^{\prime},C_{1})]=\mathsf{Q}_{L,0}^{\prime}(\bar{m})$, as in the Theorem.
For $\lambda(\bar{m})=\mathsf{P}(\bar{M}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})$, we again relabel the data, this time grouping $L,Z,M$ as $S_{t}=(L_{t},Z_{t-1},\bar{M}_{t-1})$ (and again suitably altering the meaning of $H_{t}$ to denote all variables before $A_{t}$, or $H_{A,t}$ in the original labeling). Then $\bar{M}(\bar{A}^{\mathsf{d}^{\star}})$ is simply a projection of the outcome $S_{\tau+1}$ under the longitudinal modified treatment policy $\bar{A}^{\mathsf{d}^{\star}}$. Denoting this projection as $S_{t}^{m}=\bar{M}_{t-1}$, we can write $\mathsf{P}(\bar{M}(\bar{A}^{\mathsf{d}^{\star}})=\bar{m})=\mathsf{E}[\mathbb{I}(S_{\tau+1}^{m}=\bar{m})]$, and we can again apply a similar identification strategy.
As a shorthand, let $\mathbb{I}_{\tau+1}^{\bar{m}}=\mathbb{I}(S_{\tau+1}^{m}=\bar{m})$. Let $\rho_{\tau+1}=\mathbb{I}_{\tau+1}^{\bar{m}}$, and for $t=\tau,\dots,1$ recursively define
@@
$\mathbb{I}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{P}(M_{\tau}=m_{\tau}\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau})$
$\mathbb{I}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{E}[\mathbb{I}(M_{\tau}=m_{\tau})\mid A_{\tau}=a_{\tau},H_{t}=h_{\tau}]$
@@
$\mathbb{I}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{Q}_{M,\tau}^{\star}(\underaccent{\bar}{m}_{\tau})$
@@
$\mathsf{E}[\mathbb{I}(S_{\tau}^{m}=\bar{m}_{\tau-1})\mathsf{Q}_{M,\tau}^{\star}(\underaccent{\bar}{m}_{\tau})\mid A_{\tau-1}=a_{\tau-1},H_{t}=h_{\tau-1}]$
$\mathbb{I}(S_{\tau-1}^{m}=\bar{m}_{\tau-2})\mathsf{E}[\mathbb{I}(M_{\tau-1}=m_{\tau-1})\mathsf{Q}_{M,\tau}^{\star}(\underaccent{\bar}{m}_{\tau})\mid A_{\tau-1}=a_{\tau-1},H_{t}=h_{\tau-1}]$
$\mathbb{I}(S_{\tau-1}^{m}=\bar{m}_{\tau-2})\mathsf{Q}_{M,\tau-1}^{\star}(\underaccent{\bar}{m}_{\tau-1})$
@@
$\rho_{t}(a_{t},h_{t})=\mathbb{I}(S_{t}^{m}=\bar{m}_{t-1})\mathsf{Q}_{M,t}^{\star}(\underaccent{\bar}{m}_{t})$
@@
Appendix S3 IPW identification
D´ıaz et al. (2021) gives, for a general longitudinal modified treatment policy $\bar{A}^{\mathsf{d}}$ and outcome $Y$ (without explicit mediation or intermediate confounders):
@@
Appendix S4 More on interpretation of interventional effects
@@
Interventional direct and indirect effects offer an alternative framework that overcomes this limitation. The interventional direct effect (IDE) is defined as the difference in expected outcomes when the exposure is set to $a$ and the mediator is intervened on to follow the distribution it would have under exposure $a^{\prime}$, compared to when both the exposure and mediator are set to $a^{\prime}$. Mathematically, this can be expressed as:
@@
which is not necessarily equal to $\mathsf{E}[Y(a)]-\mathsf{E}[Y(a^{\prime})]$. This is in contrast to the natural effects, since $\mathsf{E}[Y(a,M(a))]-\mathsf{E}[Y(a^{\prime},M(a^{\prime}))]$ does equal $\mathsf{E}[Y(a)]-\mathsf{E}[Y(a^{\prime})]$ since $Y(a,M(a))=Y(a)$.